Simon Companys year-end balance sheets follow. Current Yr 1 Yr Ago 2 Yrs Ago At December 31 Assets Cash Accounts receivable,Required 1 Required 2 Required 2a Required 3 Compute the return on common stockholders equity for each year. Return On CommoRequired 1 Required 2 Required 2a Required 3 Compute the price-earnings ratio for each year. (Round your answers to 2 decimalComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Assuming SimoRequired 1 Required 2 Required 2a Required 3 Compute the dividend yield for each year. (Round your answers to 2 decimal place
Answer #1

Answer:

Return on Common Stock holders Equity = Net Income / Avg. Common Stockholders Equity

Current Year = 35100 / ( 162500 + 147255 + 162500 + 95729 ) / 2 = 35100 / 283992 = 12.36%

1 Year Ago = 30541 / ( 162500 + 95729 + 162500 + 76731) / 2 = 30541 / 248730 = 12.28%

Price Earnings Ratio = Market Price / EPS

Current Year = 34 / 2.16 = 15.74

1 Year Ago = 32 / 1.88 = 17.02

Dividend Yield = Dividend per share / Market Price

Current Year = 0.28 / 34 = 0.82%

1 Year Ago = 0.14 / 32 = 0.437%